# Symmetry classification correspondence between quadratic Lindbladians and their steady states

### v1

@@ Abstract
For general coefficients, the above equation is consistent with Eq.8 only when $VC_{\text{ss}}V^{\dagger}=C_{\text{ss}}^{T}$. Then from Eq.Symmetry classification correspondence between quadratic Lindbladians and their steady states, we find that $G$ has TRS.
Crucially, this is a necessary-sufficient condition. Assume the unique steady state $G$ has TRS. We can also take the transpose of Eq. 8 and plugging the expression of $C_{\text{ss}}^{T}$ (Eq.Symmetry classification correspondence between quadratic Lindbladians and their steady states) into it. It then becomes
@@
These equations recover the constraints listed in Tab. 1 for C- symmetry with transformation $\sigma_{y}\otimes V$. Thus, the C- symmetry of $L_{\text{eff}}$ (realized by $\sigma_{y}\otimes V$) and TRS of $G$ are equivalent.
@@
Based on the above results, we can establish the correspondence between the remaining four BL classes and four of the AZ classes as the follows. When $L_{\text{eff}}$ is in the BL class BDI, it has C-- symmetry realized by $\sigma_{y}\otimes V$, and $\eta_{C}=+$. So $VV^{*}=-1$. Thus, the steady state must have TRS with $\eta_{T}=-$, i.e., must be in the AZ class AII. $L_{\text{eff}}$ in BL class CII also has C-- symmetry realized by $\sigma_{y}\otimes V$, but $\eta_{C}=-$. So the steady state must have TRS with $\eta_{T}=+$, thus must be in the AZ class AI. Similarly, the BL classes DIII and CI correspond to the AZ classes D and C, respectively.
In summary, $L_{\text{eff}}$ in BL classes AIII, BDI, DIII, CII and CI correspond to $G$ belonging to AZ classes A, AII, D, AI and C, respectively (Tab.2). The five steady state classes here are the ones that do not possess chiral symmetry. Chiral symmetry requires the appearance of both TRS and PHS. They lead to two different C- symmetries, say $U_{1}$ and $U_{2}$. They are combined to a commuting unitary symmetry: $(U_{1}U_{2}^{\dagger})L_{\text{eff}}(U_{1}U_{2}^{\dagger})^{\dagger}=-U_{1}L_{\text{eff}}^{T}U_{1}^{\dagger}=L_{\text{eff}}$. The additional unitary symmetry complicate the problem. We need to go to each symmetry sector, where the structure of $L_{\text{eff}}$ is lost. We leave these cases for future studies.
To illustrate the general applicability of our results, we provide numerical evidence from the dynamics of random Lindbladians, which can be found in EM.
Summary and discussion. In this paper, we establish a symmetry classification correspondence between quadratic fermionic Lindbladian and its steady state. We perform a BL symmetry classification on the Lindbladian and find 27 possible symmetry classes of the matrix representation $L_{\text{eff}}$. In these classes, 22 of them lead to an infinite temperature steady state. The remaining 5 classes have a one-to-one correspondence with 5 steady state AZ classes.
@@
Acknowledgements. We thank Hui Zhai, Zhong Wang, He-Ran Wang, Zijian Wang, Fei Song, Tian-Shu Deng, and Sebastian Diehl for helpful discussions.
I End Matter
I.1 Definition and construction of symmetry class
In the last 22 classes, however, $M_{l}=M_{g}^{T}$ holds. At this time, $L_{\text{eff}}$ can be transformed into block-diagonal form $UL_{\text{eff}}U^{\dagger}=\text{diag}\{-X^{T},X^{*}\}$ (see Eq.2 for definition of $X$), for $U=(\sigma_{y}+\sigma_{z})\otimes\mathbb{I}_{N}/\sqrt{2}$. So the classification of $L_{\text{eff}}$ can also be expressed by symmetries of its diagonal block $X^{T}$. We show the class of $-X^{T}\oplus X^{*}$ defined in this way in the third column of Tab.3. If we count only different classes of the diagonal block, the number of distinct classes is reduced from 22 to 7.
To see how each class is constructed, we take class $\eta_{++}$ as a representative example. In this class, $\epsilon_{Q}=+$ and $\epsilon_{QC}=+$. While the inherent Q symmetry is $\sigma_{y}\otimes\mathbb{I}_{N}$, C+ symmetry should be realized by $\mathbb{I}_{2}\otimes V_{C}$ (Tab.1). Similarly, P symmetry should be realized by $\sigma_{y}\otimes V_{P}$. Because of $\epsilon_{PC}=+$, we further demand $V_{P}V_{C}=V_{C}V_{P}$. Since $\eta_{C}=+$, we have $V_{C}V_{C}^{*}=1$.
[…]
The above construction can be carried out for all the 27 classes.
Class of $L_{\text{eff}}$ Symmetry of $L_{\text{eff}}$ $-X^{T}\oplus X^{*}$ Steady state
@@ I.1 Definition and construction of symmetry class
CII Q- and C-. $\eta_{C}=-$, $\epsilon_{QC}=+$ AZ class AI, T=+++, P=0, S=0
CI Q- and C-. $\eta_{C}=-$, $\epsilon_{QC}=-$ AZ class C, T=0, P=--, S=0
@@
$\eta_{-}$ Q- and P. $\epsilon_{PQ}=-$
@@
$\eta_{-+}$ Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=-$
$\eta_{++}$ Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=+$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=+$
$\eta_{-+}$ Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=+$ $\epsilon_{PC}=-$
@@
$\eta_{+-}$ Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=+$
$\eta_{--}$ Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=+$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=-$
$\eta_{+-}$ Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=+$ $\epsilon_{PC}=+$
@@
$\eta_{-+}$ Q-, P and C+. $\eta_{C}=-$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=-$
$\eta_{++}$ Q-, P and C+. $\eta_{C}=-$ $\epsilon_{QC}=+$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=+$
$\eta_{-+}$ Q-, P and C+. $\eta_{C}=-$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=+$ $\epsilon_{PC}=-$
@@
$\eta_{+-}$ Q-, P and C+. $\eta_{C}=-$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=+$
$\eta_{--}$ Q-, P and C+. $\eta_{C}=-$ $\epsilon_{QC}=+$, $\epsilon_{PQ}=-$ $\epsilon_{PC}=-$
$\eta_{+-}$ Q-, P and C+. $\eta_{C}=-$ $\epsilon_{QC}=-$, $\epsilon_{PQ}=+$ $\epsilon_{PC}=+$
Table 3: All possible $L_{\text{eff}}$ symmetry classes and their relation to steady state. Column one: symmetry class of $L_{\text{eff}}$, defined by column two. Column two: defining symmetries of each class. Column three: symmetry class of $-X^{T}\oplus X^{*}$, when $M_{l}=M_{g}^{T}$. Column four: Steady state properties.
I.2 Derivation of symmetry transformation reduction
@@
It is natural to introduce new free parameters $M_{+}=M_{l}+M_{g}^{T}$, $M_{-}=M_{g}^{T}-M_{l}$. The former represents total dissipation and the latter characterize fluctuations[26]. We assume the coefficients are general. Matrices $H$, $M_{+}$ and $M_{-}$ are independent and could possess some symmetries themselves.
By the form of $L_{\text{eff}}$, we should consider symmetry transformation that takes the form $U=u\otimes V$, where $u$ is a $2\times 2$ unitary and $V$ is a $N\times N$ unitary matrix. Physically, $V$ acts inside the spaces of each flavor $\mathscr{a}$ and $\mathscr{b}$. Recall that $\mathscr{a}$ and $\mathscr{b}$ are just the same fermion mode acting on different sides of the density matrix. They are subject to two physical constraints: (1) the flavors $\mathscr{a}$ and $\mathscr{b}$ should either swap or not. (2) The transformations inside $\mathscr{a}$ and $\mathscr{b}$ should be the same up to a global phase. As a result, $u$ should take the form of either $\cos\theta\mathbb{I}_{2}+i\sin\theta\sigma_{z}$ or $\cos\phi\sigma_{x}+\sin\phi\sigma_{y}$
@@
I.3 Convergence dynamics of random Lindbladians
To illustrate the validity of our results, we numerically simulate the dynamics of random Lindbladians belonging to BL classes DIII and CII with a random initial state. We show that the state converges to AZ classes D and AI at late time.
$\mathcal{I}_{\text{DIII}}$ $=\|VCV^{\dagger}-C^{T}\|_{1}/N^{2},$
$\mathcal{I}_{\text{CII}}$ $=\|VCV^{\dagger}+C^{T}-1\|_{1}/N^{2}.$ (22)

### v2

@@ V Classification of $L_{\text{eff}}$ and steady state properties
For general coefficients, the above equation is consistent with Eq.8 only when $VC_{\text{ss}}V^{\dagger}=C_{\text{ss}}^{T}$. Then from Eq.III, we find that $G$ has TRS.
Crucially, this is a necessary-sufficient condition. Assume the unique steady state $G$ has TRS. We can also take the transpose of Eq. 8 and plugging the expression of $C_{\text{ss}}^{T}$ (Eq.III) into it. It then becomes
@@
These equations recover the constraints in Tab. 1 for C- symmetry with transformation $\sigma_{y}\otimes V$. Thus, the C- symmetry of $L_{\text{eff}}$ realized by $\sigma_{y}\otimes V$ and TRS of $G$ are equivalent.
@@
Based on the above results, we can establish the correspondence between the remaining four BL classes and four of the AZ classes as the follows. When $L_{\text{eff}}$ is in the BL class BDI, it has C- symmetry realized by $\sigma_{y}\otimes V$, and $\eta_{C}=+$. So $VV^{*}=-1$. Thus, the steady state must have TRS with $\eta_{T}=$-, i.e., must be in the AZ class AII. $L_{\text{eff}}$ in BL class CII also has C- symmetry realized by $\sigma_{y}\otimes V$, but $\eta_{C}=$-. So the steady state must have TRS with $\eta_{T}=+$, thus must be in the AZ class AI. Similarly, the BL classes DIII and CI correspond to the AZ classes D and C, respectively.
In summary, $L_{\text{eff}}$ in BL classes AIII, BDI, DIII, CII and CI correspond to $G$ belonging to AZ classes A, AII, D, AI and C, respectively (Tab.2). The five steady state classes here are the ones that do not possess chiral symmetry. When steady state has chiral symmetry from the combination of TRS and PHS, $L_{\text{eff}}$ has two different C- symmetries $U_{1}=\sigma_{y}\otimes V_{T}$ and $U_{2}=\sigma_{z}\otimes V_{P}$. They are combined to a commuting unitary symmetry $(U_{1}U_{2}^{\dagger})L_{\text{eff}}(U_{1}U_{2}^{\dagger})^{\dagger}=-U_{1}L_{\text{eff}}^{T}U_{1}^{\dagger}=L_{\text{eff}}$, whose transformation is $U=U_{1}U_{2}^{\dagger}=\sigma_{x}\otimes V_{S}$. Since $V_{S}^{2}=1$, we have $U^{2}=1$ [^1] We can always achieve this goal by properly add a phase factor $e^{i\theta}$ to $U$.. So $L_{\text{eff}}$ has the block-diagonal form $L_{\text{eff}}^{+}\oplus L_{\text{eff}}^{-}$ defined by $U=\pm 1$ . For a comprehensive classification, one need to study each block $L_{\text{eff}}^{\pm}=(U\pm\mathbb{I}_{2N})L_{\text{eff}}/2$, where $V_{S}$ is involved in the classification procedure. Moreover, the inherent Q- symmetry is broken down to a relation $(\sigma_{y}\otimes\mathbb{I}_{N})L_{\text{eff}}^{+}(\sigma_{y}\otimes\mathbb{I}_{N})^{\dagger}=(L_{\text{eff}}^{-})^{\dagger}$. So the structure of $L_{\text{eff}}$ is lost. We leave these more complicated cases involving chiral symmetry for future studies.
VI Convergence dynamics of random Lindbladians
To illustrate the validity of our results, we numerically simulate the dynamics of random Lindbladians belonging to BL classes DIII and CII with a random initial state [^2] The codes for numerical calculation are available at https://github.com/maol17/2503.10763. We show that the state converges to AZ classes D and AI at late time.
$\mathcal{I}_{\text{DIII}}$ $=\|VCV^{\dagger}-C^{T}\|_{1}/N^{2},$
$\mathcal{I}_{\text{CII}}$ $=\|VCV^{\dagger}+C^{T}-1\|_{1}/N^{2}.$ (16)
VII Summary and discussion.
In this paper, we establish a symmetry classification correspondence between quadratic fermionic Lindbladian and its steady state. We perform a BL symmetry classification on the Lindbladian and find 27 possible symmetry classes. In these classes, 22 of them lead to an infinite temperature steady state. The remaining 5 classes have a one-to-one correspondence with 5 steady state AZ classes.
@@ VI Convergence dynamics of random Lindbladians
Our results potentially benefit state-preparation experiments by offering guidance for the symmetry constraints of dissipation channels. As an example, to simulate the quantum spin hall effect with, e.g., the Kane-Mele model[54, 27] (class AII, two dimensions), the Lindbladian need to be in BDI class. For the purpose of dissipative engineering, one should engineer $H$ to belong to class C, and $M_{l}$, $M_{g}$ to belong to class AII. For nearly dissipation-free systems to simulate topological dynamics, instead, TRS-breaking dissipation is the dominant channel thus should be minimized.
While applied to the $U(1)$-symmetric cases, our framework can be directly generalized to those without $U(1)$ symmetry. A generic quadratic Lindbladian takes a Bogoliubov-de Gennes (BdG) form of matrix representation $\mathscr{L}=\Psi^{\dagger}L_{\text{eff}}^{\prime}\Psi+\text{const}$, where $\Psi=(\mathscr{a},\mathscr{b},\mathscr{a}^{\dagger},\mathscr{b}^{\dagger})^{T}$. The steady state modular Hamiltonian also takes the BdG form $\hat{G}=\Phi^{\dagger}G^{\prime}\Phi$, with $\Phi=(\hat{c},\hat{c}^{\dagger})^{T}$. A similar method can be used to study the relation of $L_{\text{eff}}^{\prime}$ and $G^{\prime}$.
Furthermore, by incorporating the self-energy of modular Hamiltonian and the self-energy of Lindbladians [55], it is possible to generally study the symmetry classification correspondence for interacting fermionic systems. We leave this to future work.
VIII Acknowledgments
We thank Hui Zhai, Zhong Wang, He-Ran Wang, Zijian Wang, Fei Song, Tian-Shu Deng, and Sebastian Diehl for helpful discussions. This work is supported by the National Key R&\&D Program of China (Grant No. 2023YFA1406702) and the Start-up Research Fund of Southeast University (RF1028624190).
Class of $L_{\text{eff}}$ Symmetry of $L_{\text{eff}}$ $X^{T}\oplus X^{*}$ Steady state
@@ VIII Acknowledgments
CII Q- and C-. $\eta_{C}=$-, $\epsilon_{QC}=+$ AZ class AI, T=++, P=0, S=0
CI Q- and C-. $\eta_{C}=$-, $\epsilon_{QC}=$- AZ class C, T=0, P=-, S=0
@@
$\eta_{+}$AIII Q- and P. $\epsilon_{PQ}=+$ $\etaA\oplusη\etaA$, Q+
$\eta_{-}$AIII Q- and P. $\epsilon_{PQ}=$-
@@
$\eta_{-+}$CII Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=$-, $\epsilon_{PQ}=$- $\epsilon_{PC}=$-
$\eta_{++}$CI Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=+$, $\epsilon_{PQ}=$- $\epsilon_{PC}=+$
$\eta_{-+}$DIII Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=$-, $\epsilon_{PQ}=+$ $\epsilon_{PC}=$-
@@
$\eta_{+-}$CI Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=$-, $\epsilon_{PQ}=$- $\epsilon_{PC}=+$
$\eta_{--}$CII Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=+$, $\epsilon_{PQ}=$- $\epsilon_{PC}=$-
$\eta_{+-}$BDI Q-, P and C+. $\eta_{C}=+$ $\epsilon_{QC}=$-, $\epsilon_{PQ}=+$ $\epsilon_{PC}=+$
@@
$\eta_{-+}$BDI Q-, P and C+. $\eta_{C}=$- $\epsilon_{QC}=$-, $\epsilon_{PQ}=$- $\epsilon_{PC}=$-
$\eta_{++}$DIII Q-, P and C+. $\eta_{C}=$- $\epsilon_{QC}=+$, $\epsilon_{PQ}=$- $\epsilon_{PC}=+$
$\eta_{-+}$CI Q-, P and C+. $\eta_{C}=$- $\epsilon_{QC}=$-, $\epsilon_{PQ}=+$ $\epsilon_{PC}=$-
@@
$\eta_{+-}$DIII Q-, P and C+. $\eta_{C}=$- $\epsilon_{QC}=$-, $\epsilon_{PQ}=$- $\epsilon_{PC}=+$
$\eta_{--}$BDI Q-, P and C+. $\eta_{C}=$- $\epsilon_{QC}=+$, $\epsilon_{PQ}=$- $\epsilon_{PC}=$-
$\eta_{+-}$CII Q-, P and C+. $\eta_{C}=$- $\epsilon_{QC}=$-, $\epsilon_{PQ}=+$ $\epsilon_{PC}=+$
Table 3: All possible $L_{\text{eff}}$ symmetry classes and their relation to steady state. Column one: symmetry class of $L_{\text{eff}}$, defined by column two. Column two: defining symmetries of each class. Column three: symmetry class of $X^{T}\oplus X^{*}$, when $M_{l}=M_{g}^{T}$. Column four: Steady state properties.
IX Appendix
IX.1 Definition and construction of symmetry class
In the last 22 classes, however, $M_{l}=M_{g}^{T}$ holds. At this time, $L_{\text{eff}}$ can be transformed into block-diagonal form $UL_{\text{eff}}U^{\dagger}=\text{diag}\{-X^{T},X^{*}\}$ (see Eq.2 for definition of $X$), for $U=(\sigma_{y}+\sigma_{z})\otimes\mathbb{I}_{N}/\sqrt{2}$. So the classification of $L_{\text{eff}}$ can also be expressed by symmetries of its diagonal block $X^{T}$. We show the class of $X^{T}\oplus X^{*}$ defined in this way in the third column of Tab.3. If we count only different classes of the diagonal block, the number of distinct classes is reduced from 22 to 7.
To see how each class is constructed, we take class $\eta_{++}$BDI as a representative example. In this class, $\epsilon_{Q}=+$ and $\epsilon_{QC}=+$. While the inherent Q symmetry is $\sigma_{y}\otimes\mathbb{I}_{N}$, C+ symmetry should be realized by $\mathbb{I}_{2}\otimes V_{C}$ (Tab.1). Similarly, P symmetry should be realized by $\sigma_{y}\otimes V_{P}$. Because of $\epsilon_{PC}=+$, we further demand $V_{P}V_{C}=V_{C}V_{P}$. Since $\eta_{C}=+$, we have $V_{C}V_{C}^{*}=1$.
@@ IX.1 Definition and construction of symmetry class
The above construction can be carried out for all the 27 classes.
IX.2 Derivation of symmetry transformation reduction
@@
It is natural to introduce new free parameters $_{+}=M_{l}+M_{g}^{T}, M=M$$_{-}=M_{g}^{T}-M_{l}$. The former represents total dissipation and the latter characterize fluctuations[31]. We assume the coefficients are general. Matrices $H$, $M_{+}$ and $M_{-}$ are independent and could possess some symmetries themselves.
By the form of $L_{\text{eff}}$, we should consider symmetry transformation that takes the form $U=u\otimes V$, where $u$ is a $2\times 2$ unitary and $V$ is a $N\times N$ unitary matrix. Physically, $V$ acts inside the spaces of each flavor $\mathscr{a}$ and $\mathscr{b}$. Recall that $\mathscr{a}$ and $\mathscr{b}$ are just the same fermion mode acting on different sides of the density matrix. They are subject to two physical constraints: (1) the flavors $\mathscr{a}$ and $\mathscr{b}$ should either swap or not. (2) The transformations inside $\mathscr{a}$ and $\mathscr{b}$ should be the same up to a global phase. As a result, $u$ should take the form of either $\cos\theta\mathbb{I}_{2}+i\sin\theta\sigma_{z}$ or $\cos\phi\sigma_{x}+\sin\phi\sigma_{y}$ .